OpenStudy (eric_d):

Find value of sqrt 3 cos 15 - sin 15 @dumbcow

OpenStudy (anonymous):

0.89089

OpenStudy (dumbcow):

\[\cos \frac{x}{2} = \sqrt{\frac{1+\cos x}{2}}\] \[\sin \frac{x}{2} = \sqrt{\frac{1-\cos x}{2}}\]

OpenStudy (dumbcow):

15 is half of 30 cos 30 = sqrt3/2

OpenStudy (eric_d):

Is that a formula

OpenStudy (dumbcow):

yes plug in sqrt3/2 for cos x where x is 30 for this problem, this way you can find value for cos 15 and sin 15

OpenStudy (eric_d):

will try nw..

OpenStudy (eric_d):

 like this ? @dumbcow